Hello- Here is the question...
Each year (365 days) the fire department receives 270 emergency calls on average. In a given 7-day week, what is the probability of receiving:
a) no calls?
b) more than one call?
c) one call or less?
d) at least two calls?
I have been trying to use binomial distribution, but am confused??? Would you please help?

Answer by scott8148(6628)

binomial should work

P(c) = 270/365 ___ P(n) = 1 - (270/365)

[P(c) + P(n)]^7

a) (7C0)[P(n)]^7

b) 1 - {(7C1)[(P(c))(P(n))^6] + (7C0)[P(n)]^7

c) {(7C1)[(P(c))(P(n))^6] + (7C0)[P(n)]^7

d) same as b)